OpenStudy (anonymous):

Given the function g(x) = 8x − 2, compare and contrast g(−2) and g(4). Choose the statement that is true concerning these two values. The value of g(−2) is larger than the value of g(4). The value of g(−2) is the same as the value of g(4). The value of g(−2) is smaller than the value of g(4). The values of g(−2) and g(4) cannot be compared.

OpenStudy (kropot72):

g(-2) = -16 - 2 = -18 g(4) = 32 - 2 = 30 The value of g(-2) is smaller than the value of g(4)